Power of Tower Inc. has bonds that mature in 6½ years with a par value of RM1,000.
They pay a coupon rate of 9% with semiannual payments. If the required rate of
return on these bonds is 11% what is the bond's current value?
1
Expert's answer
2015-03-27T11:36:52-0400
N = 6½ years, F = RM1,000, if = 9% with semiannual payments, i = 11%.
The bond's current value is:
P = F*if((1 - (1 + i)^-n)/i) + F(1 + i)^-n
where:
C = F * iF = coupon payment
N = number of payments
i = market interest rate, or required yield, M = face value
P = market price of bond.
P = 1000*0.09*((1 - 1/1.11^13)/0.11) + 1000/1.11^13 = RM865.